OpenStudy (anonymous):

x^3 (third power) + 1 = -26 What is the value of X^2 (squared) +1 ?

OpenStudy (anonymous):

\[x^3 = -26-1 = -27\]\[x=-3\]\[x^2 +1 = (-3)^2 +1 = 9 +1 = 10\]